# Matrix Product State on a Quantum Computer

### v1

@@ III Turning a qMPS into canonical form
An important step before applying MPS is to turn it into canonical form. For example, Eq. 7 (8) shows how a MPS site $[A^{\delta_{i}}]_{\alpha_{i},\alpha_{i+1}}$ is in right(left)-canonical form.
@@
$\sum_{\delta_{i},\alpha_{i}}[A^{\delta_{i}}]_{\alpha_{i},\alpha_{i+1}}{\rm conj}\left([A^{\delta_{i}}]_{\alpha_{i},\alpha^{\prime}_{i+1}}\right)=\delta_{\alpha_{i+1},\alpha^{\prime}_{i+1}},$ (8)
where ${\rm conj}(A)$ indicates to take the elementwise conjugate of the tensor $A$. For classical MPS, this is done by Singular value decomposition (SVD). Unlike ref. Schuhmacher et al. (2025) that uses classical ancillary data, here we apply quantum algorhms to turn a qMPS site into canonical form, through quantum singular value decomposition (QSVD) and a quantum reshape algorithm, as shown in Fig. 2.
@@ III.1 Quantum SVD for imbalanced system
The system is divided into two subsystems, and the main idea for decomposing a state $|\psi\rangle$ is to find two unitaries $U_{A}$ and $U_{B}$ for diagonalization, such that $U_{A}\otimes U_{B}|\psi\rangle=\sum_{i}s_{i}|i\rangle|i\rangle$, as shown in Fig. 2(b) Bravo-Prieto et al. (2020). This is done through a variational algorithm optimizing $L=\sum_{q}(1-\langle\sigma_{z}^{q,A}\sigma_{z}^{q,B})/2$. We note that another variational method exists for decomposing a unitary matrix Wang et al. (2021), but in qMPS, it has to deal with a state rather than a matrix, and we have to extend the method for imbalanced systems where the number of qubits in the subsystems is unequal. Specifically for such a system, the loss function must also account for the remaining qubits, which can be updated as
$L_{\rm SVD}=\sum_{q=1}^{\min(n_{A},n_{B})}\frac{1-\langle\sigma_{z}^{q,A}\sigma_{z}^{q,B}\rangle}{2}-\sum_{q\in\mathcal{R}}(1-\langle\sigma_{z}^{q}\rangle),$ (9)
where ‘$\mathcal{R}$’ denotes the set of the remaining qubits except those in the correlated measurements. We construct a variational circuit for this purpose as shown in Fig. 2(c). Without loss of generality, we assume $n_{A}>n_{B}$. The circuit is divided into two parts, containing $n_{A}$ and $n_{B}$ qubits respectively. In each layer of the circuit, it comprises single-qubit rotations and interlacing CNOT gates. The decomposition requires that if subsystem $A$ is measured as $|i\rangle$ (in the binary representation of its $n_{A}$ qubits), then subsystem $B$ should also be in state $|i\rangle$ (in the binary representation of its $n_{B}$ qubits). This is done through the correlate measurement on the overlapped qubits of the two subsystems through the observable of $\sigma_{z}\otimes\sigma_{z}$, meanwhile the rest qubits are measured in $\sigma_{z}$ basis. Optimization is performed using standard variational algorithms, by gradient descent method. The two unitaries are obtained when the loss is optimized to zero, yielding the decomposition as
@@
where $S={\rm diag}(s_{1},s_{2},...,s_{2^{\min(n_{A},n_{B})}})$ and $U^{\prime}_{B}$ is the conjugate of $U_{B}$. As shown in Ref. Bravo-Prieto et al. (2020), the loss decays exponentially fast when the depth of the circuit grows.
@@ III.2 Quantum reshape
$L_{\rm reshape}=\sum_{q=1}^{n_{B}}(1-\langle\sigma_{z}^{q}\rangle).$ (12)
If $L_{\rm reshape}$ is optimized to 0, the output state can be determined to be $|0\rangle^{\otimes n}$ and the variational circuit produces $|\Psi_{\rm rs}\rangle$ shown in eq. 11 ideally. In our numerical simulations, we find that the variational reshape can achieve fidelities exceeding $99\%$ for circuits with 7 qubits. To prepare a right-canonical site, the corresponding data (the rows of $U^{\prime}_{B}$) can be obtained similarly.
The latter step is to keep the global state invariant by merging the rest result of QSVD (that is $SU^{\prime}_{B}$ or $U_{A}^{\dagger}S$) to the neighbouring site. We can also apply another variational algorithm to find the result of applying this operation on the neighbouring site Motta et al. (2020), but fortunately, in some cases this step is unnecessary, such as in the variational qMPS (vqMPS) method we will introduce in the next section.
@@ IV The variational MPS method on a quantum computer
Perhaps the most important application of MPS is the variational MPS (vMPS) method for finding ground energy of many-body system, which is equivalent to the DMRG method White (1992); Schollwöck (2005). With qMPS and all the components mentioned above, we can run this algorithm on quantum devices, and we denote this method by variatioanl quantum MPS (vqMPS) method. The vMPS algorithm updates the tensors site-by-site (or more than one site at a time) instead of optimizing the whole system. For site $i$, the other qMPS sites and the Hamiltonian MPO constitute its environment, or more specifically the effective Hamiltonian. This site then can be updated by the ground state of the effective Hamiltonian (see Ref. Schollwöck (2005, 2011) for detials). The vqMPS procedure closely resembles classical vMPS, by performing back-and-forth sweeps, as shown in Fig. 3.
@@
Consider updating site $i$ during a sweep from right to left, the effective hamiltonian $H_{\rm eff}^{i}$ is constructed by contracting the local tensors obtained following Eq. 6, as shown in Eq. 13.
$H_{\rm eff}^{i}=\mathcal{F}(V_{1}V_{2}\cdots V_{i-1}O_{i}V_{i+1}\cdots V_{N})$ (13)
where $\mathcal{F}$ means the contraction of all tensors with connected dimensions. We then find the ground state of $H_{\rm eff}$ through quantum algorithms like VQE or perhaps the quantum phase estimation algorithm in the future which can bring quantum advantages. The energy obtained during sweeps corresponds to the expectation value of the updated qMPS with respect to the global Hamiltonian. The number of qubits involved in this step is $12n_{\chi}+1=2\log_{2}{\chi}+$ 1. This state is then decomposed as $U_{A}^{\dagger}SU^{\prime}_{B}$ through the method mentioned above. This site can then be updated through reshaped $U^{\prime}_{B}$. Since we are going to update the $(i-1)$-th site, $U_{A}^{\dagger}S$ are no longer necessary to merge.
In practice, vqMPS can adopt a hybrid computing paradigm: Sites near boundaries with low entanglement can be efficiently updated using classical DMRG on classical computers, while central sites exhibiting strong entanglement reside on quantum devices. Consequently, this strategy leverages the advantages of both classical and quantum computing and avoids frequent quantum-classical interactions appearing in conventional VQE.
@@ V Numerical simulation
We first demonstrate the quantum SVD method for imbalanced systems. As a proof of principle, Fig. 4 shows the numerical results of decomposing random states. The circuits used are in consistency with those shown in Fig. 2(c). The result of the decomposition can be validated through the fidelity of the recovered state with the original state. We test to decompose states containing $n=3$ to $8$ qubits. For each case, we first decompose it to $n=n_{A}+n_{B}$ qubits using a variational circuit with sufficient depth, then calculate the fidelity between the original state and the state recovered from the learned parameters. Each point is averaged through 3 tests. The recovered state can reach identity fidelities for $n=3$ and $4$, and the results for more qubits are shown in Fig. 4. All results reaches fidelity higher than 96%.
@@
Figure 5: Results of numerical simulations for Heisenberg spin-chain Hamiltonian with $\Delta=0.5$, $1.0$ and $1.5$, respectively. $E_{min}$ is the ground energy found during the sweeps. The vqMPS method reaches a comparable accuracy with VQE when $n_{\chi}=1$, which requires only $32n_{\chi}+1=$ 3 qubits per VQE process. Better accuracy can be achieved when $n_{\chi}=2$ or $3$. However, for some cases using $n_{\chi}=3$ can be less accurate compared with that of $n_{\chi}=2$.
@@
where $\Delta$ is set to be $0.5$, $1.0$ and $1.5$, respectively. We test vqMPS method with bond dimensions fixed at $\chi=2,4$ or $8$, corresponding to using $n_{\chi}=1,2$ or $3$ ancillary qubits. The structure of the variational circuits are similar to that used in Fig. 2(c) but the depth of the circuit is limited to $n$ for $n$-qubit VQE. The results are shown in Fig. 5.
We performed 300 iterations and reached a good convergence for each optimization process. The maximum number of iterations is limited to control the computing cost. Since the depth of the circuits as well as the number of iterations are limited, the optimization process may not reach the ground truth. But in vqMPS, the VQE can reach a much higher ratio towards the optimal value owing to the reduced number of qubits. For VQE with less than 5 qubits, using such configurations can reach nearly optimal values, but for VQE with a large number of qubits, it requires both longer time for each step and more steps to converge. The vqMPS method can be scaled up to tens of qubits, and the time complexity is approximately linear. For all the cases, we find the vqMPS can reach a comparable accuracy compared with VQE when using merely $n_{\chi}=1$ qubit for connecting (i.e. $\chi=2$), namely each local optimization by VQE involves at most 3 qubits. The accuracy can be further improved for $n_{\chi}=2$ and $3$, using 5 or 7-qubit VQE. For some cases using $n_{\chi}=3$ can be less accurate compared with that of $n_{\chi}=2$, again reflecting the difficulty in large-scale VQE.

### v2

@@ III Turning a qMPS into canonical form
An important step before applying MPS is to turn it into canonical form. For example, Eq. 9 shows how a MPS site $A^{\delta_{i}}]_{\alpha_{i},\alpha_{i+1}}$ is in right(left)-canonical form.
@@
$\sum_{\delta_{i},\alpha_{i}}[A^{\delta_{i}}]_{\alpha_{i},\alpha_{i+1}}{\rm conj}\left([A^{\delta_{i}}]_{\alpha_{i},\alpha^{\prime}_{i+1}}\right)$ $=\delta_{\alpha_{i+1},\alpha^{\prime}_{i+1}},$
where ${\rm conj}(A)$ indicates to take the elementwise conjugate of the tensor $A$. For classical MPS, this is done by Singular value decomposition (SVD). Unlike ref. Schuhmacher2024 that uses classical ancillary data, here we apply quantum algorhms to turn a qMPS site into canonical form, through quantum singular value decomposition (QSVD) and a quantum reshape algorithm, as shown in Fig. 2.
@@ III.1 Quantum SVD for imbalanced system
The system is divided into two subsystems, and the main idea for decomposing a state $|\psi\rangle$ is to find two unitaries $U_{A}$ and $U_{B}$ for diagonalization, such that $U_{A}\otimes U_{B}|\psi\rangle=\sum_{i}s_{i}|i\rangle|i\rangle$, as shown in Fig. 2(b) Carlos2020 . This is done through a variational algorithm optimizing $L=\sum_{q}(1-\langle\sigma_{z}^{q,A}\sigma_{z}^{q,B})/2$, where $\sigma_{z}^{q,A}$ indicates the measurement on the $q$-th qubit of subsystem A. We note that another variational method exists for decomposing a unitary matrix Wang2021 , but in qMPS, it has to deal with a state rather than a matrix, and we have to extend the method for imbalanced systems where the number of qubits in the subsystems is unequal. Without loss of generality, we assume $n_{A}>n_{B}$, and the goal of the decomposition is to find the corresponding unitary such that
$U_{A}\otimes U_{B}|\psi\rangle=\sum_{i=0}^{2^{n_{B}}-1}s_{i}\left(|i\rangle\otimes|0\rangle^{\otimes{(n_{A}-n_{B})}}\right)\otimes|i\rangle.$ (10)
where $|i\rangle$ is an $n_{B}$-qubit state. The loss function must also account for the remaining qubits, which can be updated as
$L_{\rm SVD}=\sum_{q=1}^{n_{B}}\frac{1-\langle\sigma_{z}^{q,A}\sigma_{z}^{q,B}\rangle}{2}-\sum_{q=n_{B}+1}^{n_{A}}(1-\langle\sigma_{z}^{q,A}\rangle),$ (11)
We construct a variational circuit for this purpose as shown in Fig. 2(c). The circuit is divided into two parts, containing $n_{A}$ and $n_{B}$ qubits respectively. In each layer of the circuit, it comprises single-qubit rotations and interlacing CNOT gates. The decomposition requires that if subsystem $A$ is measured as $|i\rangle$ (in the binary representation of its $n_{B}$ qubits together with the latter $n_{A}-n_{B}$ qubits in state $|0\rangle$), then subsystem $B$ should also be in state $|i\rangle$ (in the binary representation of its $n_{B}$ qubits). This is done through the correlate measurement on the overlapped qubits of the two subsystems through the observable of $\sigma_{z}\otimes\sigma_{z}$, meanwhile the rest qubits are measured in $\sigma_{z}$ basis. Optimization is performed using standard variational algorithms, by gradient descent method. The two unitaries are obtained when the loss is optimized to zero, yielding the decomposition as
@@
where $S={\rm diag}(s_{1},s_{2},...,s_{2^{\min(n_{A},n_{B})}})$ and $U^{\prime}_{B}$ is the conjugate of $U_{B}$. As shown in Ref. Carlos2020 , the loss decays exponentially fast when the depth of the circuit grows. The feasibility of qSVD can be further enhanced by using a more effective ansatz or increasing parametric layers of circuit.
@@ III.2 Quantum reshape
$L_{\rm reshape}=\sum_{q=1}^{n}(1-\langle\sigma_{z}^{q}\rangle).$ (14)
If $L_{\rm reshape}$ is optimized to 0, the output state can be determined to be $|0\rangle^{\otimes n}$ and the variational circuit produces $|\Psi_{\rm rs}\rangle$ shown in eq. 13 ideally. In our numerical simulations, we find that the variational reshape can achieve fidelities exceeding $99\%$ for circuits with no more than 7 qubits. This empirical result is sufficient for our further application. To prepare a right-canonical site, the corresponding data (the rows of $U^{\prime}_{B}$) can be obtained similarly.
The latter step is to keep the global state invariant by merging the rest result of QSVD (that is $SU^{\prime}_{B}$ or $U_{A}^{\dagger}S$) to the neighbouring site. We can also apply another variational algorithm to find the result of applying this operation on the neighbouring site Motta2020 , but fortunately, in some cases this step is unnecessary, such as in the variational qMPS (vqMPS) method we will introduce in the next section.
@@ IV The variational MPS method on a quantum computer
Perhaps the most important application of MPS is the variational MPS (vMPS) method for finding ground energy of many-body system, which is equivalent to the DMRG method White1992 ; Schollwock2005 . With qMPS and all the components mentioned above, we can run this algorithm on quantum devices, and we denote this method by variatioanl quantum MPS (vqMPS) method. The vMPS algorithm updates the tensors site-by-site (or more than one site at a time) instead of optimizing the whole system. For site $i$, the other qMPS sites and the Hamiltonian MPO constitute its environment, or more specifically the effective Hamiltonian. This site then can be updated by the ground state of the effective Hamiltonian (see Ref. Schollwock2005 ; Schollwock2011 for detials). The vqMPS procedure closely resembles classical vMPS, by performing back-and-forth sweeps, as shown in Fig. 3.
@@
When updating site $i$ during a sweep from right to left, the effective Hamiltonian $H_{\rm eff}^{i}$ is constructed by contracting the local tensors obtained following Eq. 7, as shown in Eq. 15.
$H_{\rm eff}^{i}]_{\gamma^{\prime}_{i},\sigma^{\prime}_{i},\gamma^{\prime}_{i+1}}^{\gamma_{i},\sigma_{i},\gamma_{i+1}}=\sum_{\beta_{i},\beta_{i+1}}[L_{i}]^{\gamma_{i},\beta_{i},\gamma^{\prime}_{i}}[\hat{O}^{\sigma_{i},\sigma^{\prime}_{i}}]_{\beta_{i},\beta_{i+1}}[R_{i}]_{\gamma_{i+1},\beta_{i+1},\gamma^{\prime}_{i+1}}.$ (15)
where the $L$ tensor and $R$ tensor are
${}^{\gamma_{i},\beta_{i},\gamma^{\prime}_{i}}$ $=\sum_{\begin{subarray}{c}\gamma_{1},\gamma_{2},...,\gamma_{i-1}\\ \beta_{1},\beta_{2},...,\beta_{i-1}\\ \gamma^{\prime}_{1},\gamma^{\prime}_{2},...,\gamma^{\prime}_{i-1}\end{subarray}}[V_{1}]_{\gamma_{1},\beta_{1},\gamma^{\prime}_{1}}^{\gamma_{2},\beta_{2},\gamma^{\prime}_{2}}[V_{2}]_{\gamma_{2},\beta_{2},\gamma^{\prime}_{2}}^{\gamma_{3},\beta_{3},\gamma^{\prime}_{3}}\cdots[V_{i-1}]_{\gamma_{i-1},\beta_{i-1},\gamma^{\prime}_{i-1}}^{\gamma_{i},\beta_{i},\gamma^{\prime}_{i}},$ (16)
$[R_{i}]_{\gamma_{i+1},\beta_{i+1},\gamma^{\prime}_{i+1}}$ $=\sum_{\begin{subarray}{c}\gamma_{i+2},\gamma_{i+3},...,\gamma_{m+1}\\ \beta_{i+2},\beta_{i+3},...,\beta_{m+1}\\ \gamma^{\prime}_{i+2},\gamma^{\prime}_{i+3},...,\gamma^{\prime}_{m+1}\end{subarray}}[V_{i+1}]_{\gamma_{i+1},\beta_{i+1},\gamma^{\prime}_{i+1}}^{\gamma_{i+2},\beta_{i+2},\gamma^{\prime}_{i+2}}[V_{i+2}]_{\gamma_{i+2},\beta_{i+2},\gamma^{\prime}_{i+2}}^{\gamma_{i+3},\beta_{i+3},\gamma^{\prime}_{i+3}}\cdots[V_{m}]_{\gamma_{m},\beta_{m},\gamma^{\prime}_{m}}^{\gamma_{m+1},\beta_{m+1},\gamma^{\prime}_{m+1}}.$
Note that the indices are reordered, and the effective Hamiltonian $H_{\rm eff}^{i}$ is then reshaped into matrix form. Then, we find the ground state of this effective Hamiltonian, and decompose it as $U_{A}^{\dagger}SU^{\prime}_{B}$ through QSVD. This site can then be updated through reshaped $U^{\prime}_{B}$. Since we are going to update the (i1)(i-1)-th site, $U_{A}^{\dagger}S$ is no longer necessary to merge.
The ground energy of the whole system can be obtained through the sweeps resembling the classical DMRG, and our method can be scaled up to large-scale systems. For the optimization of each site, the effective Hamiltonian $H_{\rm eff}^{i}$ can be much smaller than the global Hamiltonian, and thus may mitigate the Barren Plateau by limiting the problem size. For a many-body system containing $N$ qubits, it can be represented by $N$ qMPS sites, with each site containing $2n_{\chi}+1$ qubits, and allows various means to find its ground state, such as VQE or perhaps the quantum phase estimation algorithm in the future which can bring quantum advantages. We apply VQE for the effective Hamiltonian because the main target for comparison is the full-scale VQE, using the same circuit ansatz and optimization configurations. It’s worth noting that the full-scale VQE optimizes the whole system, and is supposed to obtain better accuracy than optimizing the system site-by-site. However, it suffers from the BP phenomenon when the system reaches a certain scale. The benefit of the divide-and-conquer method is thus it can support large-scale system on limited number of qubits, while the classical DMRG theory allows to reach comparable accuracy for appropriate systems.
In practice, vqMPS can adopt a hybrid computing paradigm: Sites near boundaries with low entanglement can be efficiently updated using classical DMRG on classical computers, while central sites exhibiting strong entanglement reside on quantum devices. Consequently, this strategy leverages the advantages of both classical and quantum computing and avoids frequent quantum-classical interactions appearing in conventional VQE. A variant of VQE that applies a divide-and-conquer scheme is deepVQE Fujii2022 , which however is suitable for systems with weak correlation between subsystems. The original hybrid tensor network scheme optimizes the whole system simultaneously Yuan2021 , meaning that the optimization of each tensor is correlated with others. In our method, the optimization of each site is relatively independent once the effective Hamiltonian is given, making it more flexible.
@@ V Numerical simulation
We first demonstrate the quantum SVD method for imbalanced systems. As a proof of principle, Fig. 4 shows the numerical results of decomposing random states. The ansatz of the variational circuits used are shown in Fig. 2(c), but we use only one $R_{y}$ gate on each qubit in each variational layer for the efficiency of simulation. The result of the decomposition can be validated through the fidelity of the recovered state with the original state. We test to decompose states containing $n=3$ to $8$ qubits. For each case, we first decompose it to $n=n_{A}+n_{B}$ qubits using a variational circuit with sufficient depth, then calculate the fidelity between the original state and the state recovered from the learned parameters. Each point is averaged through 3 tests. The recovered state can reach identity fidelities for $n=3$ and $4$, and the results for more qubits are shown in Fig. 4. All simulated results reach a fidelity exceeding 96%, with fidelities above 99% for systems involving up to 7 qubits. The fidelity can be further improved by increasing the depth of the variational circuits.
@@
Figure 5: Results of numerical simulations for Heisenberg spin-chain Hamiltonian with $\Delta=0.5$, $1.0$ and $1.5$, respectively. The base-line is the result of full-scale VQE, and $E_{\rm vqMPS}-E_{VQE}$ is the difference between the ground energy found from the sweeps of vqMPS and that of full-scale VQE. The vqMPS method reaches a comparable accuracy with VQE when $n_{\chi}=1$, which requires only $2n_{\chi}+1=3$ qubits per VQE process. Better accuracy can be achieved when $n_{\chi}=2$ or 33 because all the data are below the baseline. However, for some cases using $n_{\chi}=3$ can be less accurate compared with that of $n_{\chi}=2$.
@@
where $\Delta$ is set to be $0.5$, $1.0$ and $1.5$, respectively. We test vqMPS method with bond dimensions fixed at $\chi=2,4$ or $8$, corresponding to using $n_{\chi}=1,2$ or $3$ ancillary qubits. In each case, the number of qubits involved within the optimization of a single site is $2n_{\chi}+1$, and the qSVD or qreshape can reach fidelities exceeding 99%. Note that when the full system is scaled up, the scale for optimizing each site remains the same.
We test our methods compared with the standard VQE method. The structure of the variational circuits are shown in Fig. 2(c) the depth of the circuit is limited to $n$ for $n$-qubit VQE, and for the efficiency of simulation, we use only one $R_{y}$ gate on each qubit in each variational layer. The results are shown in Fig. 5. We performed 300 iterations and reached a good convergence for each optimization process. The maximum number of iterations is limited to control the computing cost. Since the depth of the circuits as well as the number of iterations are limited, the optimization process may not reach the ground truth. But in vqMPS, the VQE can reach a much higher ratio towards the optimal value owing to the reduced number of qubits. For VQE with less than 5 qubits, using such configurations can reach nearly optimal values, but for VQE with a large number of qubits, it requires both longer time for each step and more steps to converge. The vqMPS method can be scaled up to tens of qubits, and the time complexity is approximately linear. For all the cases, we find the vqMPS can reach a comparable accuracy compared with VQE when using merely $n_{\chi}=1$ qubit for connecting (i.e. $\chi=2$), namely each local optimization by VQE involves at most 3 qubits. The accuracy can be further improved for $n_{\chi}=2$ and $3$, using 5 or 7-qubit VQE. For some cases using $n_{\chi}=3$ can be less accurate compared with that of $n_{\chi}=2$, again reflecting the difficulty in large-scale VQE.